OpenStudy (anonymous):

(-7,-6),(1,-4) write in standard form

OpenStudy (turingtest):

\[m={-4-(-6) \over 1-(-7)}={2 \over8}=1/4\] santoshf got the y's and x's mixed up...

OpenStudy (turingtest):

\[y-y_1=m(x-x_1)\] pick either point and use our m to solve for y: y-(-4)=(1/4)(x-1)=x/4-1/4 y=x/4-17/4